OpenStudy (anonymous):

using the method of substitution how do I solve 25x+5y=11 -5x=y+10

OpenStudy (mertsj):

From the second equation, y = -5x-10 Put that into the first equation in place of y

OpenStudy (anonymous):

so is it25x+5(-5x-10)=11

OpenStudy (mertsj):

\[25x+5(-5x-10)=11\]

OpenStudy (mertsj):

\[25x-25x-50=11\]

OpenStudy (mertsj):

\[0-50=11\]

OpenStudy (mertsj):

\[-50=11\]

OpenStudy (mertsj):

NO!! NOT TRUE!!! There is no solution. The lines must be parallel. Let's check the slopes and see if they are the same.

OpenStudy (mertsj):

Sure enough. The slopes are both -5 so the lines are parallel and there is no solution.

OpenStudy (anonymous):

well that is what I was wondering couldn't seem to get it was close thanks

OpenStudy (mertsj):

yw